OpenStudy (anonymous):

How to balance N2O5 +H20--->HNO3

OpenStudy (mathmale):

Would you please post this question in the Chemistry section? Thanks.

OpenStudy (tkhunny):

Solve the system of equations. a(N2O5) + b(H20) ---> 2a(N) + 5a(O) + 2b(H) + b(O) = 2a(N) + (5a+b)(O) + 2b(H) = c(HNO3) Taking the last... Matching Nitrogen 2a = c Matching Hydrogen 2b = c Matching Oxygen 5a+b = 3c With a little luck, we can find a unique solution.